OpenStudy (anonymous):

A turntable with a moment of inertia of 0.022 rotates feely at 2.7 rad/s. A circular disk of mass 370 g and diameter 30 cm, and initially not rotating, slips down a spindle and lands on the turntable.

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

This is probably an angular momentum related problem.

OpenStudy (anonymous):

sorry the question is a) Find the new angular speed. rad/s b) what is the change in kinetic energy?

OpenStudy (anonymous):

That would probably be:\[(I_a \omega_a)/(I_b+I_a)=\omega\] where I sub a is the moment of inertia of the turntable, and I sub is the moment of inertia of the circular disk. The moment of inertia of the circular disk is given by: \[Mr^2=I_b\] Where M is its total mass and r is its radius.

OpenStudy (anonymous):

I sub b is the moment of inertia of the disk*.

OpenStudy (anonymous):

I'm not sure about the change of kinetic energy of the system but that's probably: \[\Delta KE =\frac{1}{2}(I_a+I_b)\omega - \frac{1}{2} I_a \omega _a ^2\] Where omega is the new angular speed

OpenStudy (anonymous):

Thanks that really helps

OpenStudy (anonymous):

oh the omega in change in KE equation is squared.